OpenStudy (anonymous):

Help please. I got one of the 3 correct but not sure which is right. Given △BAC shown below, match the trigonometric term on the left with its corresponding ratio on the right. ______ cosine A) 12/37 ______ sine B) 35/37 ______ tangent C) 12/35

OpenStudy (anonymous):

You have to match them, put the letter in the blank. I had C, A, then B. Here is the diagram.

OpenStudy (ivettef365):

sin = opp / hyp cos = adj / hyp tan = opp / adj

OpenStudy (ivettef365):

opp = 12 adj = 35 hyp = 37

OpenStudy (ivettef365):

Can you figure it out now